# Diffusion Dynamics Models with Generative State Estimation for Cloth Manipulation

### v1

@@ III-D Model-Based Planning
We integrate our diffusion dynamics model with Model Predictive Control (MPC) for robotic cloth manipulation. Given a current cloth state $s_{t}\in\mathcal{S}$ and target state $s_{g}$, we optimize an action sequence $\{a_{t}\}_{t=0}^{T-1}$ over horizon $T$ by minimizing:
@@
where $\phi$ combines weighted MSE and chamfer distance, and $\ell$ enforces action smoothness.
@@
Our planning framework utilizes Model Predictive Path Integral (MPPI) [36] for sampling-based optimization. Actions are defined as relative end-effector displacements applied to selected cloth grasp points. To improve planning efficiency, we introduce an informed action sampling strategy and a probabilistic grasp point selection mechanism. Specifically, grasp points are selected using a temperature-controlled softmax distribution based on vertex displacements between the current and target states, while action sampling is guided by a weighted direction computed from high-displacement vertices. After each action, the robot updates its state estimate using DPM before replanning. See Appendix B-C for a detailed description of the planning algorithm and hyperparameters.
IV Experiment Setup
IV-A Task Description
We evaluate our method on challenging cloth manipulation tasks characterized by significant visual occlusion and complex physical dynamics, demonstrating the real-world performance of our diffusion-based perception and dynamics model.
Cloth folding. This task explores robotic cloth folding tasks across diverse fabrics. We employ prediction results from DPM to define target shapes, enabling accurate shape matching between the manipulated cloth and desired folding configurations. The system aims to robustly handle variations in fabric characteristics while maintaining folding accuracy. This task is more challenging than usual pushing or relocating tasks due to significant visual occlusions during the folding process, and the increased action complexity. Achieving precise folding to a specified target configuration requires both an accurate estimation and dynamic prediction of the cloth. We tested with square handkerchiefs made of three different materials. Each of these clothes has a different visual appearance and size.
T-shirt folding. This task focuses on folding a T-shirt into the target configuration. T-shirts present unique challenges due to their dual-layer structure and compliant dynamics. We evaluate our approach on four T-shirts of different sizes and physical properties. We set more challenging target states (such as diagonal fold and fold in half) that require higher motion accuracy. Incorrect actions will increase the recovery cost. Some target states also require changing the grasp contact points and performing multiple folds. Figure 3 shows all the test cloths and T-shirts with various materials and sizes used in our real-world experiments.
Figure 3: Cloth overview. We evaluate our method on different square cloths and T-shirts with varying colors and materials.
IV-B Physical Setup
IV-C Data Collection
We collect training data for learning state estimation and dynamics prediction in a simulation environment built on SAPIEN [38]. The rigid bodies, such as the robot arm, are simulated using the built-in PhysX-based simulator, while the cloth is simulated with the projective dynamics (PD) solver [4]. The two systems are coupled at the time step level by alternating updates: the PD system treats the positions and velocities of PhysX-managed objects as boundary conditions, and PhysX does the same for the PD-managed cloth. In the PD system, the cloth is modeled as a hyper-elastic thin shell. We follow Ly et al. [21] to simulate collision and friction in the PD system.
Our data collection pipeline consists of two main components: state estimation and state transitions. For state estimation, we use up to four calibrated depth cameras [45], which enable realistic stereo depth simulation, positioned at randomized angles to generate paired datasets of fused point clouds and ground-truth mesh states. The cloth is initialized by applying a random force. For state transition data, we simulate cloth trajectories by applying controlled manipulation actions. Each trajectory is generated by selecting a random vertex on the cloth and moving it to a target position (pick-and-place). We collected a comprehensive dataset of 500K examples, with action magnitudes carefully controlled between 0.02 and 0.05 units to ensure physical plausibility. Each example is an 8-step motion sequence, recording the cloth’s deformation from the initial state through the entire manipulation process. Detailed data collection procedures are provided in Appendix A-B.
IV-D Implementation Procedure
V Experiments
Figure 5: Qualitative results on state estimation. We evaluated our method on real-world deformable objects including folded cloth and T-shirts, using an RGB-D camera to capture point cloud data. Our pipeline first detects and segments the target garment, then estimates its state using DPM. Despite the challenges of significant self-occlusion, and inherent depth sensor noise, the DPM successfully reconstructs an accurate state for varied cloth and T-shirt configurations.
@@ V Experiments
$\downarrow$ MSE ($10^{-1}$) $\downarrow$ CD ($10^{-1}$) $\downarrow$ EMD ($10^{-1}$) $\downarrow$ CD ($10^{-1}$) $\downarrow$ EMD ($10^{-1}$)
@@
TABLE I: Quantitative results on state estimation. We report estimation errors in both simulated and real-world environments, with 95% confidence intervals. Lower values indicate better performance.
In this section, we investigate three key research questions:
  1. 1.
    How effectively does the Diffusion Perception Model handle self-occlusions inherent in cloth manipulation?
  2. 2.
    How does the Diffusion Dynamics Model improve dynamics prediction compared to prior approaches?
  3. 3.
    How do these enhanced perception and dynamics models translate to overall system performance?
We address these questions in three stages: evaluating state estimation accuracy (Section V-A), assessing dynamics modeling performance (Section V-B), and validating our approach through real-world cloth manipulation experiments (Section V-C).
V-A State Estimation
V-A1 Baselines
We evaluate our perception module against four baseline approaches, comprising both optimization-based and non-optimization-based prior works on cloth pose estimation, as well as ablated versions of our model.
  1. i.
    GarmentNets[7]: GarmentNets formulates the pose estimation problem of cloth as a shape completion task in the canonical space which is defined across garment instances within a category.
  2. ii.
    MEDOR[15]: MEDOR enhances the capabilities of GarmentNets by introducing test-time fine-tuning to refine mesh predictions. The approach operates in two stages: an initial forward prediction followed by an optimization process to refine the predicted mesh.
  3. iii.
    TRTM[1]: TRTM employs a template-based approach for explicit mesh reconstruction, sharing a paradigm similar to our work. However, our method diverges in both the training paradigm and model architecture, leveraging diffusion-based training and transformer-based architecture.
  4. iv.
    Transformer: To assess the effectiveness of our diffusion-based approach, we compare it against an ablated version of our model that maintains the original model architecture while removing the diffusion training.
V-A2 Results
We evaluate our method against baselines in both simulation and real-world environments using Mean Squared Error (MSE), Chamfer Distance (CD), and Earth Mover’s Distance (EMD). Quantitative results are presented in Table I. In the T-shirt object, TRTM [1] and Transformer greatly outperform GarmentNets and MEDOR [15], demonstrating that the topological information provided by the template cloth mesh significantly enhances the perception capabilities. Leveraging the cloth modeling prior during the learning process, TRTM [1] demonstrates better performance compared to Transformer. Our approach achieves further performance gains over both TRTM [1] and Transformer, highlighting the significant contributions of diffusion models to the task. We present qualitative results in Figure 5. DPM is able to accurately reconstructing the geometry of crumpled fabrics even under severe occlusion and the predicted state is also physically plausible.
V-B Dynamics Prediction Results
@@
Figure 6: Long-horizon dynamics prediction error over time. Mean Squared Error (MSE) in dynamics prediction over time under two scenarios: (a) using oracle simulation states, and (b) using DPM perception estimates, evaluated on cloths and T-shirts. Our method significantly reduces error accumulation in both cases. Error bars represent 95% confidence intervals.
V-B1 Baselines
We evaluated our diffusion dynamics models against three baseline approaches: a GNN-based method, an analytical simulator, and an ablated version of our model.
Type Method $\downarrow$ MSE $\downarrow$ CD $\downarrow$ EMD
@@ V-B1 Baselines
TABLE II: Quantitative results on long-horizon dynamics prediction. All methods use ground-truth simulator inputs to predict 36 steps forward. We report the final-frame error with a 95% confidence interval. Our DDM reduces error rates by over 10× compared to prior approaches.
  1. i.
    GNN[44]: Graph Neural Networks (GNNs) represent the most widely adopted approach for modeling dynamics. For our GNN baseline, we utilized the implementation from [44] which constructs a graph on point cloud and uses a propagation network to iteratively update the vertex and edge embedding from the encoder.
  2. ii.
    Analytical Simulator[38]: Specifically for configurations using the DPM’s output, we compared our approach against the Analytical Simulator.
  3. iii.
    Transformer: We evaluated our model against an ablated version of the diffusion dynamics module trained directly with MSE loss supervision.
For each baseline model, we analyze the mean squared error (MSE) across different timesteps on clothes and T-shirts.
V-B2 Results
Figure 7: Qualitative results on dynamics prediction. Visualization of predicted cloth configurations with point-wise L2 error, where each vertex is color-coded by its distance from the corresponding ground truth. Our diffusion-based dynamics model produces much more physically realistic deformations and significantly lower error accumulation compared to the strongest baseline, demonstrating its superior ability to model complex dynamics at high resolutions.
Our evaluation compares the proposed approach against three baselines using MSE, CD, and EMD metrics across two experimental scenarios: (1) using ground truth states from the simulator and (2) using perception states estimated by DPM. The second scenario, which includes a direct comparison with Analytical Simulator, demonstrates our method’s robustness to noisy states that typically degrade analytical simulator performance.
Error analysis over time (Figure 6) shows that DDM consistently outperforms all baselines. When using ground-truth states, GNN exhibits the lowest overall performance, except for cloth objects where it achieves comparable results to Transformer due to the simpler topology of cloth. However, for T-shirts with more complex dual-level topology, Transformer significantly outperforms GNN in final-frame MSE, highlighting the transformer architecture’s superior capability in modeling complex geometric relationships and temporal dependencies. DDM achieves the lowest MSE across all timesteps for both object types, with minimal error accumulation compared to GNN and Transformer. While Transformer’s direct MSE supervision proves effective, it fails to capture the underlying cloth state dynamics distribution. In contrast, DDM leverages diffusion models’ probabilistic nature to learn this distribution, substantially reducing temporal error accumulation. We also visualize the qualitative results in Figure 7. DDM achieves not only lower error but also produces more physically realistic dynamics predictions.
When using estimated states as inputs, where noise and physical inconsistencies are present, we include Analytical Simulator as an additional baseline for reference. While Analytical Simulator initially achieves lower error on cloth objects, its sensitivity to inconsistent inputs causes significant error accumulation, eventually leading to higher long-horizon prediction errors than DDM. This effect is even more pronounced with T-shirt objects, where the complex topology induces more sophisticated dynamics, causing Analytical Simulator’s performance to degrade more rapidly than DDM. Both GNN and Transformer exhibit even larger performance gaps under these challenging conditions. These results demonstrate that DDM achieves the best performance in realistic planning settings, where perception noise and long-horizon predictions are key challenges.
Figure 8: Qualitative results of real-world system deployment. The control framework integrates our dynamic model with Model Predictive Control (MPC) to generate and execute action sequences. We demonstrate our system’s performance across deformable objects of varying materials, appearances, and sizes. The target state is represented in the last column of each row. Each experimental sequence illustrates the progressive deformation states during folding tasks. The results validate our system’s capability to accurately manipulate diverse fabric items from arbitrary initial configurations to challenging target folding states.
@@ V-B2 Results
V-C Real World Planning Results
Comparative Analysis
We demonstrate the seamless integration of DPM and DDM within a Model Predictive Control (MPC) framework for complex cloth manipulation tasks. Our approach is benchmarked against two baselines, GNN and Transformer. Due to the unavailability of ground truth states, we employ EMD as the evaluation metric for real-world planning scenarios. Quantitative and qualitative analyses are presented in Figure 9 and Figure 8, respectively. The evaluation encompasses both single-step and multi-step prediction paradigms for cloth and T-shirt manipulation tasks.
@@ V-C Real World Planning Results
The multi-step scenarios, extending to 20 steps, further highlight our method’s efficacy in long-horizon predictions. Our approach maintains significantly reduced EMD values with a consistent downward trajectory for both cloth and T-shirt manipulation tasks. The performance gap between our method and baselines becomes increasingly pronounced over extended horizons, particularly in T-shirt manipulation where dual-layer structures introduce additional complexity. This sustained performance advantage in multi-step scenarios underscores our model’s robust capability in mitigating error accumulation while maintaining prediction accuracy across extended planning sequences.
The ablation study results presented in Table III demonstrate the critical contributions of both DPM and DDM to the system’s overall performance, with their combination yielding significantly higher success rates than either component alone.
[…]
DDM + DPM (Ours) 9/10 9/10
Transformer + DPM 3/10 2/10
GNN + DPM 6/10 5/10
DDM + Transformer 5/10 3/10
Transformer + Transformer 2/10 1/10
GNN + Transformer 5/10 4/10
TABLE III: Ablation study on planning performance. We report success rates (SR) of different variants of our method on cloth and T-shirt folding tasks.
Generalization Analysis
We demonstrate two kinds of generalization capabilities of our method.
Intra-class generalization. Our approach demonstrates strong intra-class generalization capabilities across varying cloth dimensions and material properties. We validate this by testing on square cloths ranging from 40 cm to 20 cm in side length, using a diverse set of fabrics including silk charmeuse, terry cotton, and lightweight batiste—materials that span from smooth and flowing to thick and absorbent. As shown in Figure 10, our model successfully executes precise folding trajectories across these variations, consistently achieving target configurations and confirming robust intra-class generalization.
Cross-embodiment generalization. We present qualitative cross-embodiment generalization results in Figure 11. With the embodiment-agnostic action representation, DDM enables cross-embodiment generalization by serving as a universal world model. We demonstrate this capability by successfully transferring from a parallel gripper to a five-fingered dexterous hand on folding tasks.
VI Conclusion
We introduced UniClothDiff, a unified framework that tackles key challenges in cloth manipulation by leveraging Transformer-based diffusion models for state estimation and dynamics prediction. Our approach reconstructs full cloth configurations from partial RGB-D observations and predicts long-horizon dynamics with significantly lower error than prior GNN-based methods. Integrated with model-based control, it enables precise and reliable cloth manipulation, achieving superior performance over existing approaches. Furthermore, our embodiment-agnostic action representation allows zero-shot transfer to novel robotic grippers. Through extensive experiments, we demonstrate the effectiveness of generative models in deformable object manipulation, paving the way for more robust and versatile robotic systems.
VII Limitations
@@
Figure 11: Cross-embodiment generalization results. Qualitative demonstration of our method’s performance across different robotic platforms. The table surface color is colored to yellow for improved visualization clarity.
Our approach demonstrates strong performance in cloth manipulation but has certain limitations. While transformer-based diffusion models enable accurate state estimation and long-horizon dynamics prediction, they require substantial computational resources. Additionally, achieving the intended contact point can be challenging in real-world execution when the cloth is folded, creating overlapping layers. This issue can be mitigated by integrating tactile sensing to enhance grasp precision. As our method uses a particle-based representation, it is inherently compatible with tactile feedback for more robust manipulation [2, 13], which we leave as future work.
Moreover, our method does not explicitly perform online system identification, which could improve adaptability to different object materials. However, this aspect is orthogonal to our diffusion-based dynamics model architecture. Prior work has shown that historical information can be used to infer physical parameters [44, 2], and a similar system identification phase could be readily incorporated into our framework.
Acknowledgments
This research was funded by Hillbot Inc. Hao Su is the CTO for Hillbot and receives income. The terms of this arrangement have been reviewed and approved by the University of California, San Diego in accordance with its conflict of interest policies.
Appendix A Simulation Details
A-A Simulation Setup
Detailed physical parameters for cloth simulation is presented in Table IV.
Phyical Parameter Value

### v2

@@ 3.4 Model-Based Planning
We integrate our diffusion dynamics model with Model Predictive Control (MPC) for robotic cloth manipulation. Given a current cloth state sequence $s_{t-i:t}\in\mathcal{S}$ and target state $s_{g}$, we optimize an action sequence $\{a_{t}\}_{t=0}^{T-1}$ over horizon $T$ by minimizing:
@@
where $\phi$ combines weighted mean squared error (MSE) and chamfer distance (CD), and $\ell$ enforces action smoothness. We use Model Predictive Path Integral (MPPI) [44] for sampling-based optimization. Actions are defined as relative end-effector displacements applied to a selected cloth grasp point. To improve planning efficiency, we introduce an informed action sampling strategy and a probabilistic grasp point selection mechanism. Specifically, the grasp point is selected using a temperature-controlled softmax distribution based on vertex displacements between the current and target states, while action sampling is guided by a weighted direction computed from high-displacement vertices. After each action, the robot updates its state estimate using DPM before replanning. Refer to Appendix C.4 for details on the planning algorithm and hyperparameters.
4 Experiments
We investigate three key research questions: (1) How effectively does the Diffusion Perception Model handle self-occlusions inherent in cloth manipulation? (2) How does the Diffusion Dynamics Model improve dynamics prediction compared to prior approaches? (3) How do these enhanced perception and dynamics models translate to overall system performance? We study these questions by evaluating state estimation accuracy (Section 4.2), assessing dynamics modeling performance (Section 4.3), and real-world deployment across two system setups(Section 4.4).
4.1 Experiments Setup
We evaluate our method in both simulation and real-world environments. Specifically, we use SAPIEN [45] as the simulation platform for data collection and training, and demonstrate effective sim-to-real transfer in the real-world setting. Additional details of the pipeline implementation and experimental setup are provided in Appendix B.
4.2 State Estimation Results
Figure 2: Qualitative results on state estimation. Row (a) shows raw point cloud of the work space. Row (b) shows the segmented point clouds of real-world clothes, all of which are highly crumpled. Row (c) shows the predicted cloth states.
Baselines. We compare our perception module against four baselines: GarmentNets [3] which formulates cloth pose estimation problem as a shape completion task in the canonical space; MEDOR [4] which improves GarmentNets by introducing test-time fine-tuning for mesh refinement; TRTM [5] which employs a template-based approach for explicit mesh reconstruction; and Transformer, an ablated version of our model that retains the original architecture but without diffusion training. These baselines cover both optimization-based and non-optimization-based prior works on cloth pose estimation, along with ablation studies for our model.
Results. We evaluate our method and baselines in both simulation and the real world using MSE, CD, and Earth Mover’s Distance (EMD). The results are presented in Table 1. In the T-shirt setting, TRTM and Transformer greatly outperform GarmentNets and MEDOR, demonstrating that the topological information provided by the template cloth mesh significantly enhances the perception capabilities. Leveraging the cloth modeling prior during the learning process, TRTM demonstrates better performance compared to Transformer. Our approach achieves further performance gains over both TRTM and Transformer, highlighting the significant contributions of diffusion models to the task. We provide qualitative results in Figure 2.
@@ 4.2 State Estimation Results
MSE $\downarrow$ ($10^{-1}$) CD $\downarrow$ ($10^{-1}$) EMD $\downarrow$ ($10^{-1}$) CD $\downarrow$ ($10^{-1}$) EMD $\downarrow$ ($10^{-1}$)
@@
Table 1: Quantitative results on state estimation. We report estimation errors in both simulated and real-world scenarios, with 95% confidence intervals. Lower values indicate better performance.
4.3 Dynamics Prediction Results
Baselines. We evaluated our diffusion dynamics models against three baseline approaches: a GNN-based method [6] which is the most widely adopted approach for modeling dynamics; an Analytical Simulator specifically for configurations using the DPM’s output; and an ablated version of our model with dynamics module trained directly with MSE loss supervision termed Transformer. For each baseline model, we analyze the MSE across different timesteps on clothes and T-shirts.
Results. Our evaluation compares the proposed approach against three baselines using MSE across two experimental scenarios: (1) using ground truth states from the simulator and (2) using perception states estimated by DPM. The second scenario, which includes a direct comparison with Analytical Simulator, demonstrates the robustness of our method to noisy states that typically degrade the performance of the analytical simulator.
Error analysis over time in Figure 3 shows that DDM consistently outperforms all baselines. GNN exhibits the weakest performance, particularly for complex objects like T-shirts. Transformer improves over GNN by leveraging transformer architectures, but still suffers from error accumulation. In contrast, DDM achieves the lowest MSE across all timesteps with minimal temporal error accumulation, benefiting from the diffusion model’s expressive distribution modeling. Qualitative results in Figure 4 further highlight the physical plausibility of DDM’s predictions.
@@ 4.3 Dynamics Prediction Results
Figure 3: Long-horizon dynamics prediction error over time. MSE in dynamics prediction over time under two scenarios: (a) using oracle simulation states, and (b) using DPM perception estimates, evaluated on clothes and T-shirts. Error bars represent 95% confidence intervals.
When using estimated states with perception noise, we introduce Analytical Simulator as an additional baseline. Although Analytical Simulator initially achieves low error on cloth objects, it is highly sensitive to inconsistent inputs, leading to rapid error accumulation and worse long-horizon performance than DDM. This degradation is even more pronounced for T-shirt objects due to their complex topology. Under these realistic conditions, both GNN and Transformer exhibit even larger gaps, demonstrating that DDM provides the most robust and accurate dynamics predictions for planning with noisy observations. We provide additional quantitative results in Appendix A.1.
Figure 4: Qualitative results on dynamics prediction. Visualization of predicted clothes configurations with each vertex color-coded by its point-wise L2 error from ground truth.
4.4 Real World Planning Results
Method Cloth T-shirt Long-sleeve
Self Ext. Comb. Self Ext. Comb. Self Ext. Comb.
GNN 6/10 4/10 3/10 1/10 2/10 2/10 2/10 2/10 0/10
Ours 9/10 8/10 6/10 9/10 7/10 6/10 7/10 6/10 4/10
Table 2: Quantitative results of real-world manipulation. We repeat each scenario for 10 trials, with randomized initial and target states.
[…]
DDM + DPM (Ours) 9/10 8/10
Transformer + DPM 3/10 5/10
GNN + DPM 6/10 1/10
DDM + Transformer 5/10 3/10
Transformer + Transformer 2/10 1/10
GNN + Transformer 5/10 0/10
Table 3: Success rates of system variants with different combinations of dynamics and perception modules.
Comparative Analysis. We demonstrate the seamless integration of DPM and DDM within an MPC framework for complex cloth folding tasks. Our approach is benchmarked against GNN as a dynamics module. We categorize our tasks into folding and unfolding scenarios, all of which involve long-horizon challenges and require multi-step prediction, with three distinct occlusion types: self-occlusion, external occlusion by other objects (e.g., a robotic arm), and combined occlusion, which poses challenges for accurate perception in cloth manipulation. We report the success rates (SR) of the quantitative results in Table 2. A trial is deemed successful if the geometric metric (EMD) is below a certain threshold (Appendix A.2). Our method consistently outperforms GNN across all occlusion scenarios. In simpler tasks, such as folding clothes, our model achieves an improvement of approximately 30% points in SR. When manipulating more challenging objects, such as a dual-layer T-shirt where GNN struggles to accurately model dynamics, our approach achieves up to a 50% increase in SR. Qualitative results are shown in Figure 5, illustrating challenging initial and target configurations with severe self-occlusion, representing substantially more difficult setups than those considered in many prior works [4, 6, 17, 25]. By leveraging an embodiment-agnostic action space design, our approach enables effective cross-embodiment transfer. We demonstrate a successful transfer of the embodiment from a parallel gripper to a dexterous hand in Figure 6.
Figure 5: Qualitative results on real-world cloth manipulation. We evaluate our system on three types of garment: square clothes, long-sleeve shirts and T-shirts. For each garment, the first row corresponds to folding and the second row to unfolding.
Figure 6: Cross-embodiment generalization results.
Ablation. Ablation study results (Table 3) demonstrate the critical contributions of both DPM and DDM to the system’s overall performance, with their combination yielding significantly higher success rates than either component alone. For objects with simpler topology, such as cloth, accurate perception is most critical since the dynamics are relatively straightforward and easy to model. However, for objects with more complex topologies, having an accurate dynamics model becomes equally important for effective planning.
5 Conclusion
We introduce UniClothDiff, a unified framework that tackles key challenges in state estimation and dynamics prediction in cloth manipulation with Transformer-based diffusion models. Our approach reconstructs full cloth configurations from partial RGB-D observations and predicts long-horizon dynamics with significantly lower error than prior GNN-based methods. Integrated with model-based control, it enables cloth manipulation in various scenarios, significantly outperforming existing approaches. Through extensive experiments, we demonstrate the potential of generative models for deformable object manipulation, paving the way for more robust and versatile robotic systems.
6 Limitations
One limitation of our method is the substantial computational cost associated with training large transformer-based diffusion models. Second, our experiments primarily focus on cloth manipulation; extending the framework to contact-rich rigid body tasks is a promising future direction. Given the generality of the model design, we expect feasible adaptation with suitable training data. Additionally, our model currently lacks explicit uncertainty estimates. Incorporating uncertainty quantification into perception and dynamics models, and combining them with control methods with theoretical guarantees, could improve robustness in safety-critical scenarios.
7 Acknowledgments
This research was funded by Hillbot Inc. Hao Su is the CTO for Hillbot and receives income. The terms of this arrangement have been reviewed and approved by the University of California, San Diego in accordance with its conflict of interest policies. We thank Xuhui Kang for his help with video production and editing and Zhan Ling for his helpful discussions.
Appendix A Additional Results
A.1 Dynamics Prediction
Quantitative Results
We provide additional quantitative results of forward dynamics prediction for both ground truth state and perception noisy input in Table 4 and Table 5. Our dynamics model consistently outperforms all baselines across both evaluation scenarios. When using ground-truth states from the simulator as input, DDM achieves approximately 10× lower error than the best-performing baseline, highlighting the superiority of diffusion models in capturing long-horizon dynamics. When using estimated states from DPM, which introduces additional noise, DDM still achieves 2× lower error than all baselines, demonstrating that the diffusion-based training paradigm significantly enhances noise tolerance through its expressive data distribution modeling capacity.
Type Method MSE $\downarrow$ CD $\downarrow$ EMD $\downarrow$
@@ A.1 Dynamics Prediction
Table 4: Quantitative results on dynamics prediction with ground truth input. Errors represent a 95% confidence interval.
Type Method MSE $\downarrow$ CD $\downarrow$ EMD $\downarrow$
($10^{-3}$) ($10^{-2}$) ($10^{-2}$)
T-shirt GNN 6.36 $\pm$ 1.30 8.88 $\pm$ 1.12 8.29 $\pm$ 1.94
Transformer 4.18 $\pm$ 0.73 4.26 $\pm$ 0.51 7.93 $\pm$ 0.70
DDM 0.55 $\pm$ 0.27 1.49 $\pm$ 0.13 3.22 $\pm$ 0.47
Cloth GNN 2.17 $\pm$ 1.44 5.02 $\pm$ 0.90 7.31 $\pm$ 4.65
Transformer 1.30 $\pm$ 0.65 2.27 $\pm$ 0.46 7.06 $\pm$ 2.08
DDM 0.66 $\pm$ 0.45 2.12 $\pm$ 0.54 5.51 $\pm$ 1.03
Table 5: Quantitative results on dynamics prediction with perception input. Errors represent a 95% confidence interval.
Qualitative Results
We present additional qualitative results for dynamics prediction in Figure 7 and Figure 8. Each row represents a predicted dynamics sequence. The results demonstrate the physical plausibility of the generated outputs.
A.2 Real-world Planning
Additional Quantitative Results
We present quantitative results using the EMD metric, which measures the distance from the initial state to the target state in real-world planning scenarios in Figure 9.
@@ A.2 Real-world Planning
The multi-step scenarios, extending to 20 steps, further highlight our method’s efficacy in long-horizon predictions. Our approach maintains significantly reduced EMD values with a consistent downward trajectory for both cloth and T-shirt manipulation tasks. The performance gap between our method and baselines becomes increasingly pronounced over extended horizons, particularly in T-shirt manipulation, where dual-layer structures introduce additional complexity. This sustained performance advantage in multi-step scenarios underscores our model’s robust capability in mitigating error accumulation while maintaining prediction accuracy across extended planning sequences.
Figure 10: Success rate under different threshold.
We present a comprehensive breakdown of the success rates in real-world long-horizon settings under different thresholds in Figure 10, where our method consistently outperforms all baselines.
Additional Qualitative Results
Accordingly, we also provide additional qualitative results for single-step and multi-step scenarios on clothes and T-shirts in Figure 11. The results validate our system’s capability to accurately manipulate diverse fabric items from arbitrary initial configurations to challenging target folding states.
Figure 11: Qualitative results of real-world system deployment. The target state is represented in the last column of each row. Each experimental sequence illustrates the progressive deformation states during folding tasks. The first two rows correspond to single-step scenarios, while the last two represent multi-step scenarios.
Additional Intra-class Generalization Results
We provide additional qualitative results for intra-class generalization on square cloth object in Figure 12 with sizes ranging from 20 cm to 40 cm. Our model successfully executes precise folding trajectories across these variations, consistently achieving target configurations and confirming robust intra-class generalization.
@@
A.3 Simulation Planning Results
We present more qualitative results in Figure 14 in the simulation environment on planning. We design four simple tasks in simulation for system validation as visualized in Figure 13.
(a) Lift cloth
(b) Fold cloth
(c) Rotate cloth
(d) Pull cloth
Appendix B Experiment Setup
B.1 Task Description
Figure 15: Coverage of different clothes in our experiments.
Figure 17: Cloth overview. We evaluate our method on different single-layer clothes as well as dual-layer T-shirts and long-sleeve shirts with varying colors and materials.
We evaluate our method on challenging cloth folding tasks characterized by significant visual occlusion and complex physical dynamics, demonstrating the real-world performance of our diffusion-based perception and dynamics model.
Square Cloth folding and unfolding. This task explores robotic cloth folding and unfolding tasks across diverse fabrics. We employ prediction results from DPM to define target shapes, enabling accurate shape matching between the manipulated cloth and desired folding configurations. The system aims to robustly handle variations in fabric characteristics while maintaining folding accuracy. This task is more challenging than usual pushing or relocating tasks due to significant visual occlusions during the folding process, and the increased action complexity. Achieving precise folding to a specified target configuration requires both an accurate estimation and the dynamic prediction of the cloth. We tested with square handkerchiefs made of three different materials. Each of these clothes has a different visual appearance and size.
Garment folding and unfolding. This task focuses on folding or unfolding a T-shirt or a long-sleeve top into the target configuration. Such garments present unique challenges due to their dual-layer structure and compliant dynamics. We evaluate our approach on garments of different sizes and physical properties. We set more challenging target states (such as diagonal fold and fold in half) that require higher motion accuracy. Incorrect actions will increase the recovery cost. Some target states also require changing the grasp contact points and performing multiple folds. Figure 17 shows all the test cloths and garments with various materials and sizes used in our real-world experiments.
We visualize the distribution of all the clothes and garments in our experiments in Figure 17
B.2 Physical Setup
Appendix C Implementation Details
C.1 Data Collection
We collect training data for learning state estimation and dynamics prediction in a simulation environment built on SAPIEN [45]. The rigid bodies, such as the robot arm, are simulated using the built-in PhysX-based simulator, while the cloth is simulated with the projective dynamics (PD) solver [46]. The two systems are coupled at the time step level by alternating updates: the PD system treats the positions and velocities of PhysX-managed objects as boundary conditions, and PhysX does the same for the PD-managed cloth. In the PD system, the cloth is modeled as a hyper-elastic thin shell. We follow Ly et al. [47] to simulate collision and friction in the PD system. We provide detailed physical parameters for cloth simulation in Table 6.
Physical Parameter Value